# Multi-Scale Spatial-Temporal Hypergraph Network with Lead-Lag Structures for Stock Time Series Forecasting

### v1

@@ Abstract
Time series forecasting occurs in a range of financial applications providing essential decision-making support to investors, regulatory institutions, and analysts. Unlike multivariate time series from other domains, stock time series exhibit industry correlation. Exploiting this kind of correlation can improve forecasting accuracy. However, existing methods based on hypergraphs can only capture industry correlation relatively superficially. These methods face two key limitations: they do not fully consider inter-industry lead-lag interactions, and they do not model multi-scale information within and among industries. This study proposes the Hermes framework for stock time series forecasting that aims to improve the exploitation of industry correlation by eliminating these limitations. The framework integrates moving aggregation and multi-scale fusion modules in a hypergraph network. Specifically, to more flexibly capture the lead-lag relationships among industries, Hermes proposes a hyperedge-based moving aggregation module. This module incorporates a sliding window and utilizes dynamic temporal aggregation operations to consider lead-lag dependencies among industries. Additionally, to effectively model multi-scale information, Hermes employs cross-scale, edge-to-edge message passing to integrate information from different scales while maintaining the consistency of each scale. Experimental results on multiple real-world stock datasets show that Hermes outperforms existing state-of-the-art methods in both efficiency and accuracy.
@@ 1 Introduction
Time series data, which organizes information chronologically by timestamps (Wu et al., 2025a; Qiu et al., 2025e; Liu et al., 2025a; Hu et al., 2025; Dai et al., 2024b; Sun et al., 2025a), has become increasingly prevalent with ongoing digitalization in various fields (Qiu et al., 2025d; Wu et al., 2025b; Liu et al., 2025b; Sun et al., 2025b). Among them, stock time series are a prominent example. A prominent example is stock time series data, which typically comprises records from multiple stocks, each containing key financial indicators such as opening prices, closing prices, and more (Fan & Shen, 2024; Li et al., 2024). Stock time series forecasting (STSF) is important in financial applications, providing crucial decision-making support for investors, regulatory institutions, and analysts, helping them make data-driven decisions in complex and ever-changing market environments (Rather et al., 2017; Faloutsos et al., 2018; Jacob & Shasha, 1999; Zhang et al., 2014; Zheng et al., 2023).
Most existing studies treat stocks in stock time series independently, thus disregarding the interconnected nature of market dynamics and contradicting the intrinsic interdependencies observed in real-world financial systems (Diebold & Yılmaz, 2014). In reality, stocks are often interrelated, and there are rich signals in the relationships between them (Nobi et al., 2014). For example, with the breakthroughs of Artificial Intelligence technology, companies like Apple, Microsoft, and Alphabet that actively invest and innovate in this domain are expected to lead in the market due to their technological advantages. It is one of the important factors for their stock prices to show a steady upward trend in the overall stock market111www.morningstar.co.uk/uk/news/258865/tech-and-communication-stocks-drove-us-market-gains-in-2024.aspx—see Figure 1(a). At the same time, the soaring power demand, driven by electrification, decarbonization, and the advent of artificial intelligence, led to a concurrent rise in the stock prices of companies in the energy sector [^2] https://finance.yahoo.com/news/american-electric-power-aep-among-121752028.html?—see Figure 1(a). Such industry correlations have gradually attracted more attention. Especially within the same industry, different stocks often exhibit correlations, reflecting the common economic environment, market fluctuations, policy changes, and other factors that impact the stocks.
@@
Challenge 1: Current methods fail to fully consider inter-industry lead-lag interactions. In financial markets, there are causal relationships and time dependencies among industries, and these relationships are highly diverse and complex. As shown in Figure 1(b), the technology industry has flourished with breakthroughs in key technologies and market innovations, achieving significant progress. Technological innovations from technology companies have injected new vitality into the development of renewable energy, and energy companies have leveraged technologies such as big data, AI, and the Internet of Things to enhance the efficiency of energy production and management2, thereby driving up stock prices (Abid et al., 2025; Olanrewaju et al., 2024; Rafiei et al., 2025). Therefore, in the modeling process, it is important to capture lead-lag correlations among industries more accurately, so as to better reflect the dynamic interactions and dependencies in financial markets, ultimately improving the accuracy and practicality of the forecasts.
@@
In this study, we address the two challenges by proposing a general stock time series forecasting framework, Hermes, which integrates hyperedge-based moving aggregation and hyperedge-based multi-scale fusion modules into a spatial-temporal hypergraph network. These integrations enhance the precision and stability of stock forecasting. First, to more flexibly model lead-lag relationships among industries, we integrate a hyperedge-based moving aggregation module into the hypergraph network. This module incorporates a sliding window and utilizes dynamic temporal aggregation operations to consider lead-lag dependencies among industries. This approach aims to capture a market’s complex structure more precisely. Second, to address the issue that existing methods fail to effectively model multi-scale information, we decompose raw data at multiple scales and integrate a multi-scale fusion module into the hypergraph network. This module employs cross-scale, edge-to-edge message passing to integrate information from different scales while preserving the consistency of each scale, thereby boosting both the accuracy and stability of forecasts. Experimental results on multiple real-world stock datasets show that Hermes outperforms existing state-of-the-art methods in both efficiency and accuracy. Our contributions are summarized as follows.
  • To solve the STSF problem, we propose a general spatial-temporal hypergraph network framework, Hermes, which learns more accurate and adaptive forecasting models by considering both lead-lag and multi-scale industry correlations.
  • We design a hyperedge-based multi-scale fusion module. This module utilizes a cross-scale, egde-to-edge message-passing to integrate information from different scales while maintaining the consistency of each scale.
@@ 2.1 Deep Learning Stock Time Series Forecasting
With the rise of deep learning (Huang et al., 2025; Fu et al., 2025), notable progress has occurred in stock forecasting. This is mainly because deep neural networks can model complex nonlinear patterns. Some studies use recurrent neural networks (RNNs) (Nelson et al., 2017; Qin et al., 2017; Akita et al., 2016; Rahman et al., 2019) to capture the historical development in individual stock prices and predict their short-term trends. In addition, due to the significant progress of Multi-Layer Perceptrons (MLPs) architecture in the general time series field (Zeng et al., 2023; Lin et al., 2025; 2024). Increasingly may proposals target stock forecasting performance by improving the MLPs architecture. For example, a series of studies (Fan & Shen, 2024; Lazcano et al., 2024; Tashakkori et al., 2024) enhance the learning ability of simple MLPs structures by utilizing the MLP-Mixer backbone, thereby improving forecasting accuracy. Next, the Transformer model has become increasingly popular due to the ability of its self-attention mechanism to capture long-term dependencies. For example, Master (Li et al., 2024) proposes a novel stock transformer for stock price forecasting to effectively capture stock correlation. While these deep learning methods have advanced stock forecasting, the complexity of stock data and the need to account for intricate relationships among stocks have led researchers to explore new methods based on hypergraphs.
@@ 2.2 Hypergraph-based Stock Time Series Forecasting
Traditional graphs have limitations when it comes to modeling higher-order multivariate relationships, as they can only represent relations between pairs of nodes. To address this issue, Hypergraph Neural Networks (HGNNs) (Feng et al., 2019c) were introduced. Hypergraphs extend the concept of simple graphs to enable the capture of relationships among multiple stocks (Chen et al., 2020). Hypergraphs have gained increasing attention and application across various fields. Hyperedge in a hypergraph represents a set of vertices, making them suitable for modeling non-pairwise relationships (Luo et al., 2014; Sawhney et al., 2020; Huynh et al., 2023). For example, HGAM (Li et al., 2022) is a hypergraph-based reinforcement learning method for stock portfolio selection. THGNN (Xiang et al., 2022) is a temporal and heterogeneous graph neural network model that aims to learn dynamic relationships by using two-stage attention mechanisms. STHAN-SR (Sawhney et al., 2021) enhances corporate relevance based on Wiki data and uses hypergraph convolution to propagate information from higher-order neighbors. ESTIMATE (Huynh et al., 2023), employs hypergraphs to capture non-pairwise correlations, utilizing temporal generative filters to identify individual patterns for each stock. Unlike the above methods, we propose a new hypergraph-based stock time series forecasting model that specifically takes into account the multi-scale nature of stock data, enabling it to capture the complexity of market dynamics more comprehensively. Furthermore, our approach places particular emphasis on the lead-lag relationships between financial industries, where the changes in one industry may precede or lag behind those in another. By capturing these key relationships more fully, we aim to improve forecasting accuracy.
@@ 3 Methodology
In financial markets, a stock time series ${\bm{\mathsfit{X}}}=\left\{{\bm{X}}_{1},\ldots,{\bm{X}}_{N}\right\}$ records the historical data of $N$ stocks, where ${\bm{X}}_{i}\in\mathbb{R}^{T\times F}$ is data of one stock with a lookback window of $T$ time steps and with $F$ technical indicators at each time step. In stock time series, the indicators may include opening price, closing price, highest price, lowest price, and trading volume. Following existing studies (Huynh et al., 2023; Fan & Shen, 2024), we input a stock time series with multiple indicators, denoted as ${\bm{\mathsfit{X}}}\in\mathbb{R}^{N\times T\times F}$, and aim to predict the closing price $p_{i}^{t}$ at time step $t$. We define the 1-day return ratio as $r_{i}^{t}=\frac{p_{i}^{t}-p_{i-1}^{t}}{p_{i-1}^{t}}$. With $\theta$ representing the model parameters, the process can be expressed as follows.
${\bm{\mathsfit{X}}}\in\mathbb{R}^{N\times T\times F}\stackrel{{\scriptstyle\theta}}{{\rightarrow}}p\in\mathbb{R}^{N\times 1}\rightarrow r\in\mathbb{R}^{N\times 1}$ (1)
@@ 3.1 Framework Overview
In this section, we introduce the framework of our proposed spatiotemporal hypergraph neural network that considers multiple time scales and models the lead-lag relationships between industries. Our framework takes historical financial sequences as input, first samples the time series into several sequences of different granularities according to different time scales, and maps the feature dimension to the hidden dimension to consider the high-dimensional correlations between time series segments under different scales. Subsequently, we consider constructing relationships between agents at each scale, specifically considering constructing hyperedges for agents in the same industry, and modeling the lead-lag relationships between industries in the corresponding hypergraphs. Then, we model the interactions between different scales, construct hyperedges for industries of different scales, and adaptively learn the correlations between them under the corresponding multi-scale hypergraph, thereby playing a role in fusing scales. Finally, we predict the corresponding indicators for each agent for the next day through a prediction head based on a multilayer perceptron.
@@ 3.2 Multi-Scale Feature Extraction
Specifically, for the financial input time series ${\bm{\mathsfit{X}}}\in\mathbb{R}^{N\times T\times F}$, which denotes the records of $F$ key indicators for $N$ market agents of $T$ timestamps in the past. To decouple the intricate multi-resolution information, we use convolutional layers of different sizes to adaptively down-sample the time series into sequences of different scales:
@@
Feature Extraction. Subsequently, for time series of different scales, we enhance their representations at two aspects. First, we adopt the Causal-Mixing technique to process the temporal dimensional. The Causal-Mixing technique ensures causal relationships rather than full connectivity, and ensure that time series points can only see previous points. It consists of a group of MLP layers:
$\text{Causal-Mixing}^{i}=\{\text{MLP}^{i,1},\text{MLP}^{i,2},\cdots,\text{MLP}^{i,T^{i}}\},~~\text{MLP}^{i,j}=\mathbb{R}^{j}\to\mathbb{R}$ (3)
$\mathbb{X}^{i}=\text{Causal-Mixing}^{i}({\bm{\mathsfit{X}}}^{i}),$ (4)
the Causal-Mixing is applied along the temporal dimensional, which maps $X^{i}_{1:j}$ through $\text{MLP}^{i,j}$ to $\mathbb{X}^{i}_{j}$. The new representation $\mathbb{X}\in\mathbb{R}^{N\times T^{i}\times F}$ keeps the original shape.
@@
$\mathcal{X}^{i}=\mathbb{X}^{i}\cdot\text{W}^{i},$ (5)
@@ 3.3 Adaptive Spatial-Temporal Hypergraph Construction
where $\mathcal{X}^{i}$ is the representation of $i$-the scales and works as the nodes of the $\mathcal{G}^{i}$, $E^{i}\in\mathbb{R}^{K\times T^{i}\times d}$ is the hyperedge constructed through the stock-industry relationship matrix $H$:
$\text{scores}^{i}=\text{Linear}^{i}(\mathcal{X}^{i})\in\mathbb{R}^{N\times 1},~~\text{Linear}^{i}=\mathbb{R}^{T^{i}\times d}\to\mathbb{R},$ (7)
$H^{i}=H\odot\text{scores}^{i}+(1-H)\odot(-\infty),~~\mathcal{H}^{i}=\text{SoftMax}(H^{i},\text{dim}=0),~~E^{i}$ $=(\mathcal{H}^{i})^{T}\cdot\mathcal{X}^{i},$ (8)
where the matrix $H$ is first processed with a learnable renovation to adapt to the current representation $\mathcal{X}^{i}$. $H^{i}$ is the intermediate result obtained through endowing the $H_{l,m}=1$ parts with learnable $\text{scores}^{i}$ and $H_{l,m}=0$ parts with -$\infty$ to ensure zeros in the SoftMax step. Then the $\mathcal{H}^{i}\in\mathbb{R}^{N\times K}$ with Continuous Weights is obtained, and the hyperedges $E^{i}\in\mathbb{R}^{K\times T^{i}\times d}$ of industries are generated through information integration from all the related agents. We obtain all the Spatial-Temporal hypergraphs $\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots\mathcal{G}^{S}\}$ for each scale as above mentioned.

### v2

# Hermes: A Multi-Scale Spatial-Temporal Hypergraph Network for Stock Time Series Forecasting

@@ Abstract
Time series forecasting occurs in a range of financial applications providing essential decision-making support to investors, regulatory institutions, and analysts. Unlike multivariate time series from other domains, stock time series exhibit industry correlation. Exploiting this kind of correlation can improve forecasting accuracy. However, existing methods based on hypergraphs can only capture industry correlation relatively superficially. These methods face two key limitations: they do not fully consider inter-industry lead-lag interactions, and they do not model multi-scale information within and among industries. This study proposes the Hermes framework for stock time series forecasting that aims to improve the exploitation of industry correlation by addressing these limitations. The framework integrates moving aggregation and multi-scale fusion modules in a hypergraph network. Specifically, to more flexibly capture the lead-lag relationships among industries, Hermes proposes a hyperedge-based moving aggregation module. This module incorporates a sliding window and utilizes dynamic temporal aggregation operations to consider lead-lag dependencies among industries. Additionally, to effectively model multi-scale information, Hermes employs cross-scale, edge-to-edge message passing to integrate information from different scales while maintaining the consistency of each scale. Experimental results on multiple real-world stock datasets show that Hermes outperforms existing state-of-the-art methods.
@@ 1 Introduction
Stock time series are typically composed of data from multiple stocks, each of which includes multiple key indicators such as stock opening prices, closing prices, etc. Fan and Shen (2024); Li et al. (2024). Stock time series forecasting (STSF) is important in financial applications, providing crucial decision-making support for investors, regulatory institutions, and analysts, helping them make data-driven decisions in complex and ever-changing market environments Rather et al. (2017); Faloutsos et al. (2018); Jacob and Shasha (1999); Zheng et al. (2023).
Most existing studies treat stocks in stock time series independently, thus disregarding the interconnected nature of market dynamics and contradicting the intrinsic interdependencies observed in real-world financial systems Diebold and Yılmaz (2014). In reality, stocks are often interrelated, and there are rich signals in the relationships between them Nobi et al. (2014). For example, with the breakthroughs of Artificial Intelligence technology, companies like Apple, Microsoft, and Alphabet that actively invest and innovate in this domain are expected to lead in the market due to their technological advantages. It is one of the important factors for their stock prices to show a steady upward trend in the overall stock market111www.morningstar.co.uk/uk/news/258865/tech-and-communication-stocks-drove-us-market
-gains-in-2024.aspx
—see Figure 1(a). At the same time, the soaring power demand, driven by electrification, decarbonization, and the advent of artificial intelligence, led to a concurrent rise in the stock prices of companies in the energy sector [^2] https://finance.yahoo.com/news/american-electric-power-aep-among-121752028.html?—see Figure 1(a). Such industry correlations have gradually attracted more attention. Especially within the same industry, different stocks often exhibit correlations, reflecting the common economic environment, market fluctuations, policy changes, and other factors that impact the stocks.
@@
Challenge 1: Current methods fail to fully consider inter-industry lead-lag interactions. In financial markets, there are time dependencies among industries, and these relationships are highly diverse and complex. As shown in Figure 1(b), the technology industry has flourished with breakthroughs in key technologies and market innovations, achieving significant progress. Technological innovations from technology companies have injected new vitality into the development of renewable energy, and energy companies have leveraged technologies such as big data, AI, and the Internet of Things to enhance the efficiency of energy production and management2, thereby driving up stock prices Abid et al. (2025); Olanrewaju et al. (2024); Rafiei et al. (2025). Therefore, in the modeling process, it is important to capture lead-lag correlations among industries more accurately, so as to better reflect the dynamic interactions and dependencies in financial markets, ultimately improving the accuracy and practicality of the forecasts.
@@
In this study, we address the two challenges by proposing a general stock time series forecasting framework, Hermes, which integrates hyperedge-based moving aggregation and hyperedge-based multi-scale fusion modules into a spatial-temporal hypergraph network. These integrations enhance the precision and stability of stock forecasting. First, to more flexibly model lead-lag relationships among industries, we integrate a hyperedge-based moving aggregation module into the hypergraph network. This module incorporates a sliding window and utilizes dynamic temporal aggregation operations to consider lead-lag dependencies among industries. This approach aims to capture a market’s complex structure more precisely. Second, to address the issue that existing methods fail to effectively model multi-scale information, we decompose raw data at multiple scales and integrate a multi-scale fusion module into the hypergraph network. This module employs cross-scale, edge-to-edge message passing to integrate information from different scales while preserving the consistency of each scale, thereby boosting both the accuracy and stability of forecasts. Experimental results on multiple real-world stock datasets show that Hermes outperforms existing state-of-the-art methods.
Our contributions are summarized as follows.
  • To solve the STSF problem, we propose a general spatial-temporal hypergraph network framework, Hermes, which learns more accurate and adaptive forecasting models by considering lead-lag and multi-scale industry correlations.
@@
  • We design a hyperedge-based multi-scale fusion module. This module utilizes a cross-scale, edge-to-edge message-passing to integrate information from different scales while maintaining the consistency of each scale.
@@ 2.1 Deep Learning Stock Time Series Forecasting
With the rise of deep learning, notable progress has occurred in stock forecasting. This is mainly because deep neural networks can model complex nonlinear patterns. Some studies use recurrent neural networks (RNNs) Nelson et al. (2017); Qin et al. (2017); Akita et al. (2016); Rahman et al. (2019) to capture the historical development in individual stock prices and predict their short-term trends. In addition, due to the significant progress of Multi-Layer Perceptrons (MLPs) architecture in the general time series field Zeng et al. (2023); Lin et al. (2025, 2024). Increasingly many proposals target stock forecasting performance by improving the MLPs architecture. For example, a series of studies Fan and Shen (2024); Lazcano et al. (2024); Tashakkori et al. (2024) enhance the learning ability of simple MLPs structures by utilizing the MLP-Mixer backbone, thereby improving forecasting accuracy. Next, the Transformer model has become increasingly popular due to the ability of its self-attention mechanism to capture long-term dependencies. For example, Master Li et al. (2024) proposes a novel stock transformer for stock price forecasting to effectively capture stock correlation. While these deep learning methods have advanced stock forecasting, the complexity of stock data and the need to account for intricate relationships among stocks have led researchers to explore new methods based on hypergraphs.
@@ 2.2 Hypergraph-based Stock Time Series Forecasting
Traditional graphs have limitations when it comes to modeling higher-order multivariate relationships, as they can only represent relations between pairs of nodes. To address this issue, Hypergraph Neural Networks (HGNNs) Feng et al. (2019c) were introduced. Hypergraphs extend the concept of simple graphs to enable the capture of relationships among multiple stocks Chen et al. (2020). Hypergraphs have gained increasing attention and application across various fields. Hyperedge in a hypergraph represents a set of vertices, making them suitable for modeling non-pairwise relationships Luo et al. (2014); Sawhney et al. (2020); Huynh et al. (2023). For example, HGAM Li et al. (2022) is a hypergraph-based reinforcement learning method for stock portfolio selection. THGNN Xiang et al. (2022) is a temporal and heterogeneous graph neural network model that aims to learn dynamic relationships by using two-stage attention mechanisms. STHAN-SR Sawhney et al. (2021) enhances corporate relevance based on Wiki data and uses hypergraph convolution to propagate information from higher-order neighbors. ESTIMATE Huynh et al. (2023), employs hypergraphs to capture non-pairwise correlations, utilizing temporal generative filters to identify individual patterns for each stock. Unlike the above methods, we propose a new hypergraph-based stock time series forecasting model that specifically takes into account the multi-scale nature of stock data, enabling it to capture the complexity of market dynamics more comprehensively. Furthermore, our approach places particular emphasis on the lead-lag relationships between financial industries, where the changes in one industry may precede or lag behind those in another. By capturing these key relationships more fully, we aim to improve forecasting accuracy.
@@ 3 Methodology
In financial markets, a stock time series ${\bm{\mathsfit{X}}}=\left\{{\bm{X}}_{1},\ldots,{\bm{X}}_{N}\right\}$ records the historical data of $N$ stocks, where ${\bm{X}}_{i}\in\mathbb{R}^{T\times F}$ is data of one stock with a lookback window of $T$ time steps and with $F$ technical indicators at each time step. In stock time series, the indicators may include opening price, closing price, highest price, lowest price, and trading volume. Following existing studies Huynh et al. (2023); Fan and Shen (2024), we input a stock time series with multiple indicators, denoted as ${\bm{\mathsfit{X}}}\in\mathbb{R}^{N\times T\times F}$, and aim to predict the closing price $p_{i}^{t}$ at time step $t$. We define the 1-day return ratio as $r_{i}^{t}=\frac{p_{i}^{t}-p_{i}^{t-1}}{p_{i}^{t-1}}$. With $\theta$ representing the model parameters, the process can be expressed as follows.
${\bm{\mathsfit{X}}}\in\mathbb{R}^{N\times T\times F}\stackrel{{\scriptstyle\theta}}{{\rightarrow}}p\in\mathbb{R}^{N\times 1}\rightarrow r\in\mathbb{R}^{N\times 1}.$ (1)
@@ 3.1 Framework Overview
In this section, we introduce the framework of our proposed spatiotemporal hypergraph neural network that considers multiple time scales and models the lead-lag relationships between industries—see Figure 2. Our framework takes historical financial sequences as input, first samples the time series into several sequences of different granularities according to different time scales, and maps the feature dimension to the hidden dimension to consider the high-dimensional correlations between time series segments under different scales. Subsequently, we consider constructing relationships between stocks at each scale, specifically considering constructing hyperedges for stocks in the same industry, and modeling the lead-lag relationships between industries in the corresponding hypergraphs. Then, we model the interactions between different scales, construct hyperedges for industries of different scales, and adaptively learn the correlations between them under the corresponding multi-scale hypergraph, thereby playing a role in fusing scales. Finally, we predict the corresponding indicators for each stock for the next day through a prediction head based on a multilayer perceptron.
@@ 3.2 Multi-Scale Feature Extraction
Specifically, for the financial input time series ${\bm{\mathsfit{X}}}\in\mathbb{R}^{N\times T\times F}$, which denotes the records of $F$ key indicators for $N$ stocks of $T$ timestamps in the past. To decouple the intricate multi-resolution information, we use convolutional layers of different sizes to adaptively down-sample the time series into sequences of different scales:
@@
Feature Extraction. Subsequently, for time series of different scales, we enhance their representations at two aspects. First, we adopt the Causal-Mixing technique to process the temporal dimension. The Causal-Mixing technique ensures causal relationships rather than full connectivity, and ensure that time series points can only see previous points. It consists of a group of MLP layers:
$\text{Causal-Mixing}^{i}=\{\text{MLP}^{i,1},\text{MLP}^{i,2},\cdots,\text{MLP}^{i,T^{i}}\},$ (3)
$\text{MLP}^{i,j}=\mathbb{R}^{j}\to\mathbb{R},$ (4)
$\mathbb{X}^{i}=\text{Causal-Mixing}^{i}({\bm{\mathsfit{X}}}^{i}).$ (5)
The Causal-Mixing is applied along the temporal dimension, which maps $X^{i}_{1:j}$ through $\text{MLP}^{i,j}$ to $\mathbb{X}^{i}_{j}$. The new representation $\mathbb{X}\in\mathbb{R}^{N\times T^{i}\times F}$ keeps the original shape.
@@
$\mathcal{X}^{i}=\mathbb{X}^{i}\cdot W^{i},$ (6)
@@ 3.3 Adaptive Spatial-Temporal Hypergraph Construction
where $\mathcal{X}^{i}$ is the representation of $i$-th scale and works as the nodes of the $\mathcal{G}^{i}$, $E^{i}\in\mathbb{R}^{K\times T^{i}\times d}$ is the hyperedge constructed through the stock-industry relationship matrix $H$:
$\text{scores}^{i}=\text{Linear}^{i}(\mathcal{X}^{i})\in\mathbb{R}^{N\times 1},$ (8)
$\text{Linear}^{i}=\mathbb{R}^{T^{i}\times d}\to\mathbb{R},$ (9)
$H^{i}=H\odot\text{scores}^{i}+(1-H)\odot(-\infty),$ (10)
$\mathcal{H}^{i}=\text{SoftMax}(H^{i},\text{dim}=0),$ (11)
$E^{i}=(\mathcal{H}^{i})^{T}\cdot\mathcal{X}^{i},$ (12)
where the matrix $H$ is first processed with a learnable renovation to adapt to the current representation $\mathcal{X}^{i}$. $H^{i}$ is the intermediate result obtained through endowing the $H_{l,m}=1$ parts with learnable $\text{scores}^{i}$ and $H_{l,m}=0$ parts with -$\infty$ to ensure zeros in the SoftMax step. Then the $\mathcal{H}^{i}\in\mathbb{R}^{N\times K}$ with Continuous Weights is obtained, and the hyperedges $E^{i}\in\mathbb{R}^{K\times T^{i}\times d}$ of industries are generated through information integration from all the related stocks. We obtain all the Spatial-Temporal hypergraphs $\{\mathcal{G}^{1},\mathcal{G}^{2},\cdots\mathcal{G}^{S}\}$ for each scale as above mentioned.